OpenStudy (anonymous):

MEDAL AND FAN What are the vertical asymptotes of the function f(x) = the quantity of 4 x plus 8, all over x squared plus 3 x minus 4

OpenStudy (anonymous):

i think x=-1 and x=2

OpenStudy (anonymous):

does anyone know if i am right

OpenStudy (usukidoll):

\[f(x) = \frac{4x+8}{x^2+3x-4}\]

OpenStudy (usukidoll):

is this your function?

OpenStudy (usukidoll):

I'm gonna assume it is. Anyway , Vertical asymptotes are vertical lines which correspond to the zeroes of the denominator of a rational function. so we have to factor \[x^2+3x-4\]

OpenStudy (usukidoll):

\[(x+1)(x-4) = 0\] solve for x x+1 = 0 and x-4 = 0 only x = -1 is correct so try solving x-4=0

OpenStudy (anonymous):

4

OpenStudy (usukidoll):

yup x = 4, and x = -1 are our vertical asymptotes good job :)